# Deep Proxy Causal Learning and its Application to Confounded Bandit Policy Evaluation

### v4

@@ 1 Introduction
In causal learning, we aim to estimate the effect of our actions on the world. For example, we may be interested in measuring the impact of flight ticket prices on sales [Wright, 1928, Blundell et al., 2012], or the effect of grade retention on cognitive development [Fruehwirth et al., 2016]. We refer to our action as a treatment, which results in a particular outcome. It is often impossible to determine the effect of treatment on outcome from observational data alone, since the observed joint distribution of treatment and outcome can depend on a common confounder which influences both, and which might not be observed. In our example on sales of plane tickets given a price, the two might even be positively correlated in some circumstances, such as the simultaneous increase in sales and prices during the holiday season. This does not mean that raising the price causes sales to increase. In this context, people’s desire to fly is a confounder, since it affects both the number of ticket sales and the prices people are willing to accept. Thus, we need to correct the bias caused by the confounder.
@@
One common assumption to cope with confounding bias is to assume no unobserved confounders exist [Imbens, 2003], or more generally, the ignorable treatment assignment assumption [Rosenbaum and Rubin, 1983], which states that the treatment assignment is independent of the potential outcomes caused by the treatment, given the background data available. Although a number of methods are proposed based on this assumption [Hill, 2011, Johansson et al., 2016, Yao et al., 2018], it can be too restrictive, since it is often difficult to determine how the confounder affects treatment assignments and outcomes.
@@
A less restrictive assumption is that we have access to proxy variables, which contain relevant side information on the confounder. In the flight tickets example, we can use the number of views of the ticket reservation page as a proxy variable, which reflects peoples’ desire for flights. Note that if we can completely recover the confounder from proxy variables, the ignorable treatment assignment assumption can be satisfied. Motivated by this, Lee et al. [2018] and Louizos et al. [2017] aim to recover the distribution of confounders from proxy variables using modern machine learning techniques such as generative adversarial networks [Goodfellow et al., 2014] or variational auto-encoders (VAE) Kingma and Welling [2014]. Although these methods exhibit powerful empirical performance, there is little theory that guarantees the correct recovery of the causal effects.
@@
Kuroki and Pearl [2014] first considered the necessary conditions on proxy variables to provably recover the underlying causal effect via direct recovery of the hidden confounder. This work was in turn generalized by Miao et al. [2018a]. In their work, it is shown that two types of proxy variables are sufficient to recover the true causal effects without explicitly recovering the confounder. One is an outcome-inducing proxy, which correlates with confounders and causes the outcome, and the other is a treatment-inducing proxy which is caused by confounders and correlates with the treatment. In the flight ticket example, we can use the number of views of the ticket reservation page as the outcome-inducing proxy, and the cost of fuel as the treatment-inducing proxy. Given these proxy variables, Miao et al. [2018a] show that the true causal effect can be recovered by solving a Fredholm integral equation, which is referred to as the proxy causal learning (PCL) problem. The PCL problem is known to have an interesting relation to the causal inference with multiple treatments, which uses a subset of treatments as proxy variables in PCL [Wang and Blei, 2021, Miao et al., 2020].
@@
Although the PCL problem has a solid theoretical grounding, the question of how to estimate the causal effect remains a practical challenge, in particular when we consider nonlinear causal relationships or high-dimensional treatments. In Miao et al. [2018a], the treatment and outcome are assumed to be categorical variables. In a follow-up study, Miao et al. [2018b] show that we can learn a linear causal effect by a method of moments. Deaner [2018] models the causal effect as a linear combination of nonlinear basis functions, which is learned by solving two stage regression. These two methods are extended by Mastouri and Zhu et al. [Mastouri et al., 2021], who learn the causal effect in a predefined reproducing kernel Hilbert space (RKHS). We provide an overview of the PCL problem and the two-stage regression in Section 2. Although these methods enjoy desirable theoretical properties, the flexibility of the model is limited, since all existing work uses pre-specified features.
@@
In this paper, we propose a novel Deep Feature Proxy Variable (DFPV) method, which is the first work to apply neural networks to the PCL problem. The technique we employ builds on earlier work in instrumental variable (IV) regression, which is a related causal inference setting to PCL. A range of deep learning methods has recently been introduced for IV regression [Hartford et al., 2017, Xu et al., 2021, Bennett et al., 2019]. We propose to adopt the Deep Feature Instrumental Variable method Xu et al. [2021], which learns deep adaptive features within a two-stage regression framework. Details of DFPV are given in Section 3. In Section 4, we empirically show that DFPV outperforms other PCL methods in several examples. We further apply PCL methods to the off-policy evaluation problem in a confounded bandit setting, which aims to estimate the average reward of a new policy given data with confounding bias. We discuss the setting in Section 3, and show the superiority of DFPV in experiments in Section 4.
@@ 2.1 Problem Setting for Proxy Causal Learning
Assumption 1 (Structural Assumption [Deaner, 2018, Mastouri et al., 2021]).
@@
Assumption 2 (Completeness Assumption on Confounder [Deaner, 2018, Mastouri et al., 2021]).
@@
Figure 1 shows the causal graph describing these relationships. In our setting, we assume that there is no observable confounder, though this may be easily included [Mastouri et al., 2021, Tchetgen et al., 2020] as presented in Appendix D. Here, the bidirectional arrows mean that we allow both directions or even a common ancestor variable. Given these assumptions, it is shown that the structural function can be expressed using a solution of an integral equation.
Proposition 1 (Miao et al. [2018a]).
@@
For completeness, we present a proof in Appendix B (Lemma 2 and Corollary 1), which is due to Miao et al. [2018b] and Deaner [2018]. Following Miao et al. [2018b], we call $h^{*}$ the bridge function. From Proposition 1, we can see that the estimation of the structural function reduces to the estimation of the bridge function, since once we obtain the bridge function, the structural function directly follows from (2).
@@ 2.2 Two-stage Regression Scheme for Proximal Causal Learning with Fixed Features
To obtain the bridge function, we need to solve the functional equation (1). However, directly solving (1) in a rich function space can be ill-posed (see discussion in Nashed and Wahba [1974]). To address this, recent works [Deaner, 2018, Mastouri et al., 2021] minimize the following regularized loss $\mathcal{L}_{\mathrm{PV}}$ to obtain an estimate of the bridge function $\hat{h}$:
@@
Deaner [2018] and Mastouri and Zhu et al. [Mastouri et al., 2021]solve the minimization problem (3) by successively solving two-stage regression problems. They model the bridge function as
@@
where $\bm{u}$ is a learnable weight vector, $\bm{\psi}_{A}(a),\bm{\psi}_{W}(w)$ are vectors of fixed basis functions, and $\otimes$ is a Kronecker product, defined as $\bm{a}\otimes\bm{b}=\mathrm{vec}(\bm{a}\bm{b}^{\top})$ for any finite dimensional vectors $\bm{a},\bm{b}$ [^1] Throughout this paper, we assume the number of basis functions to be finite. Mastouri and Zhu et al. [Mastouri et al., 2021] consider an infinite number of basis function in a reproducing kernel Hilbert space, and use the definitions of the inner and Kronecker products for that space. An estimate $\hat{\bm{u}}$ is obtained by solving the successive regression problems. In Stage 1, we estimate the conditional expectation $\mathbb{E}_{W|A=a,Z=z}\left[\bm{\psi}_{W}(W)\right]$ as a function of $a,z$. In Stage 2, we substitute the model (4) into the inner conditional expectation in $\mathcal{L}_{\mathrm{PV}}$,
@@
When fixed feature dictionaries are used, this two-stage regression benefits from strong theoretical guarantees [Deaner, 2018, Mastouri et al., 2021]. The use of pre-specified feature maps, limits the scope and flexibility of the method, however, especially if the treatment and proxies are high dimensional (e.g. images or text), and the causal relations are nonlinear. To overcome this, we propose to use adaptive features, expressed by neural nets, as described in the next section.
@@ 3 Deep Feature Proxy Causal Learning
In this section, we develop the DFPV algorithm, which learns adaptive features modeled by neural nets using a technique similar to Xu et al. [2021]. As in Mastouri and Zhu et al. [Mastouri et al., 2021], we assume that we do not necessarily have access to observations from the joint distribution of $(A,Y,Z,W)$. Instead, we are given $m$ observations of $(A,Z,W)$ for Stage 1 and $n$ observations of $(A,Z,Y)$ for Stage 2. We denote the stage 1 observations by $(a_{i},z_{i},w_{i})$ and the stage 2 observations by $(\tilde{a}_{i},\tilde{z}_{i},\tilde{y}_{i})$. If observations of $(A,Y,Z,W)$ are given for both stages, we can evaluate the out-of-sample loss of Stage 1 using Stage 2 data and vice versa, and these losses can be used for hyper-parameter tuning of $\lambda_{1},\lambda_{2}$ (Appendix A). We first introduce two-stage regression with adaptive feature maps and then describe the detailed learning procedure of DFPV.
@@ 3.1 Two-stage regression with adaptive features
As in the existing work [Deaner, 2018, Mastouri et al., 2021], we learn $\mathbb{E}_{W|a,z}\left[\bm{\psi}_{\theta_{W}}(w)\right]$ in Stage 1 and $h(a,w)$ in Stage 2, but in addition to the weights $\bm{u}$ and $\bm{V}$, we also learn the parameters of the feature maps. Specifically, in Stage 1, we learn $\bm{V}$ and parameters $\theta_{A(1)},\theta_{Z}$ by minimizing the following empirical loss:
@@
Again, the loss ^$\hat{\mathcal{L}}_{2}$ is an empirical estimate of $\mathcal{L}_{2}$ in (6) with adaptive feature maps. Although the expression of ^$\hat{\mathcal{L}}_{2}$ does not explicitly contain $\theta_{W}$, it implicitly depends on $\theta_{W}$ as $(\hat{\bm{V}},\hat{\theta}_{A(1)},\hat{\theta}_{Z})$ is the solution of a minimization problem involving $\theta_{W}$. This implicit dependency makes it challenging to update $\theta_{W}$, as we cannot directly obtain its gradient. One possible solution is to use the implicit gradient method [Lorraine et al., 2020], but this approach might be too computationally expensive. Instead, we use the method proposed in Xu et al. [2021], in which we ignore the dependency of $\theta_{W}$ on parameters $\hat{\theta}_{A(1)},\hat{\theta}_{Z}$, and compute the gradient via the closed-form solution of $\hat{V}$.
@@ 3.2 Deep Feature Proxy Variable Method
The complete procedure is presented in Algorithm 1. Note that we may use any sophisticated gradient-based learning method to optimize, such as Adam [Kingma and Ba, 2015]. As reported in Xu et al. [2021], we observe that the learning procedure is stabilized by running several gradient descent steps on the stage 1 parameters $(\theta_{A(1)},\theta_{Z})$ before updating the stage 2 features $(\theta_{A(2)},\theta_{W})$. Furthermore, we may use mini-batch updates, which sample subsets of the data at the beginning of each iteration and only use these subsamples to update the parameters.
@@ 3.3 Consistency of DFPV
In this section, we show that the DFPV method yields a consistent estimate of the bridge function. Our approach differs from the consistency result of Deaner [2018] and Mastouri and Zhu et al. [Mastouri et al., 2021], which depend on the assumption that the true bridge function lies in a specific functional space (such as an RKHS [Mastouri et al., 2021]) or satisfies certain smoothness properties [Deaner, 2018]. Since our features are adaptive, however, we instead build our result on a Rademacher complexity argument [Mohri et al., 2012], which holds for a wide variety of hypothesis spaces. Furthermore, we do not assume that our hypothesis space contains the true bridge function, which makes this theory of independent interest.
@@
Assumption 3 (Completeness Assumption on Outcome-Inducing Proxy [Deaner, 2018, Mastouri et al., 2021]).
@@
We formally restate Proposition 2 in Appendix C (Theorem 3), in which the definitions of hypothesis spaces and constants can be found. From this, we can see that if we correctly identify the hypothesis space ($\kappa_{1},\kappa_{2}=0$) and have vanishing Rademacher complexities (^$\hat{\mathfrak{R}}_{S_{1}}(\mathcal{H}_{1}),\hat{\mathfrak{R}}_{S_{2}}(\mathcal{H}_{2})\to 0$ in probability), then the estimated bridge function converges in probability to the true one. Note that the Rademacher complexity of certain neural networks is known; e.g., ReLU two-layer networks are known to have the complexity of order $O(\sqrt{1/n})$ [Neyshabur et al., 2019]. However, it might be not straightforward to derive the Rademacher complexities of the specific classes of neural networks we use, because we are employing the outer product of two neural networks in (5) and (6), whose Rademacher complexity remains a challenging open problem. The analysis in Proposition 2 is similar to Xu et al. [2021], with two important distinctions: first, the target $f_{\mathrm{struct}}$ is estimated using the bridge function $h$ marginalized over one of its arguments $W$, but PCL learns the bridge function itself. Second, our result is more general and includes the misspecified case, which was not treated in the earlier work.

### v5

@@ 1 Introduction
In causal learning, we aim to estimate the effect of our actions on the world. For example, we may be interested in measuring the impact of flight ticket prices on sales [34, 2], or the effect of grade retention on cognitive development [4]. We refer to our action as a treatment, which results in a particular outcome. It is often impossible to determine the effect of treatment on outcome from observational data alone, since the observed joint distribution of treatment and outcome can depend on a common confounder which influences both, and which might not be observed. In our example on sales of plane tickets given a price, the two might even be positively correlated in some circumstances, such as the simultaneous increase in sales and prices during the holiday season. This does not mean that raising the price causes sales to increase. In this context, people’s desire to fly is a confounder, since it affects both the number of ticket sales and the prices people are willing to accept. Thus, we need to correct the bias caused by the confounder.
@@
One common assumption to cope with confounding bias is to assume no unobserved confounders exist [8], or more generally, the ignorable treatment assignment assumption [28], which states that the treatment assignment is independent of the potential outcomes caused by the treatment, given the background data available. Although a number of methods are proposed based on this assumption [7, 9, 36], it can be too restrictive, since it is often difficult to determine how the confounder affects treatment assignments and outcomes.
@@
A less restrictive assumption is that we have access to proxy variables, which contain relevant side information on the confounder. In the flight tickets example, we can use the number of views of the ticket reservation page as a proxy variable, which reflects peoples’ desire for flights. Note that if we can completely recover the confounder from proxy variables, the ignorable treatment assignment assumption can be satisfied. Motivated by this, Lee et al. [14] and Louizos et al. [17] aim to recover the distribution of confounders from proxy variables using modern machine learning techniques such as generative adversarial networks [5] or variational auto-encoders (VAE) [11]. Although these methods exhibit powerful empirical performance, there is little theory that guarantees the correct recovery of the causal effects.
@@
Kuroki and Pearl [13] first considered the necessary conditions on proxy variables to provably recover the underlying causal effect via direct recovery of the hidden confounder. This work was in turn generalized by Miao et al. [20]. In their work, it is shown that two types of proxy variables are sufficient to recover the true causal effects without explicitly recovering the confounder. One is an outcome-inducing proxy, which correlates with confounders and causes the outcome, and the other is a treatment-inducing proxy which is caused by confounders and correlates with the treatment. In the flight ticket example, we can use the number of views of the ticket reservation page as the outcome-inducing proxy, and the cost of fuel as the treatment-inducing proxy. Given these proxy variables, Miao et al. [20] show that the true causal effect can be recovered by solving a Fredholm integral equation, which is referred to as the proxy causal learning (PCL) problem. The PCL problem is known to have an interesting relation to the causal inference with multiple treatments, which uses a subset of treatments as proxy variables in PCL [33, 22].
@@
Although the PCL problem has a solid theoretical grounding, the question of how to estimate the causal effect remains a practical challenge, in particular when we consider nonlinear causal relationships or high-dimensional treatments. In Miao et al. [20], the treatment and outcome are assumed to be categorical variables. In a follow-up study, Miao et al. [21] show that we can learn a linear causal effect by a method of moments. Deaner [3] models the causal effect as a linear combination of nonlinear basis functions, which is learned by solving two stage regression. These two methods are extended by Mastouri and Zhu et al. [18], who learn the causal effect in a predefined reproducing kernel Hilbert space (RKHS). We provide an overview of the PCL problem and the two-stage regression in Section 2. Although these methods enjoy desirable theoretical properties, the flexibility of the model is limited, since all existing work uses pre-specified features.
@@
In this paper, we propose a novel Deep Feature Proxy Variable (DFPV) method, which is the first work to apply neural networks to the PCL problem. The technique we employ builds on earlier work in instrumental variable (IV) regression, which is a related causal inference setting to PCL. A range of deep learning methods has recently been introduced for IV regression [6, 35, 1]. We propose to adopt the Deep Feature Instrumental Variable method [35], which learns deep adaptive features within a two-stage regression framework. Details of DFPV are given in Section 3. In Section 4, we empirically show that DFPV outperforms other PCL methods in several examples. We further apply PCL methods to the off-policy evaluation problem in a confounded bandit setting, which aims to estimate the average reward of a new policy given data with confounding bias. We discuss the setting in Section 3, and show the superiority of DFPV in experiments in Section 4.
@@ 2.1 Problem Setting for Proxy Causal Learning
Assumption 1 (Structural Assumption [3, 18]).
@@
Assumption 2 (Completeness Assumption on Confounder [3, 18]).
@@
Figure 1 shows the causal graph describing these relationships. In our setting, we assume that there is no observable confounder, though this may be easily included [18, 31] as presented in Appendix D. Here, the bidirectional arrows mean that we allow both directions or even a common ancestor variable. Given these assumptions, it is shown that the structural function can be expressed using a solution of an integral equation.
Proposition 1 (Miao et al. [20]).
@@
For completeness, we present a proof in Appendix B (Lemma 2 and Corollary 1), which is due to Miao et al. [21] and Deaner [3]. Following Miao et al. [21], we call $h^{*}$ the bridge function. From Proposition 1, we can see that the estimation of the structural function reduces to the estimation of the bridge function, since once we obtain the bridge function, the structural function directly follows from (2).
@@ 2.2 Two-stage Regression Scheme for Proximal Causal Learning with Fixed Features
To obtain the bridge function, we need to solve the functional equation (1). However, directly solving (1) in a rich function space can be ill-posed (see discussion in Nashed and Wahba [25]). To address this, recent works [3, 18] minimize the following regularized loss $\mathcal{L}_{\mathrm{PV}}$ to obtain an estimate of the bridge function $\hat{h}$:
@@
Deaner [3] and Mastouri and Zhu et al. [18]solve the minimization problem (3) by successively solving two-stage regression problems. They model the bridge function as
@@
where $\bm{u}$ is a learnable weight vector, $\bm{\psi}_{A}(a),\bm{\psi}_{W}(w)$ are vectors of fixed basis functions, and $\otimes$ is a Kronecker product, defined as $\bm{a}\otimes\bm{b}=\mathrm{vec}(\bm{a}\bm{b}^{\top})$ for any finite dimensional vectors $\bm{a},\bm{b}$ [^1] Throughout this paper, we assume the number of basis functions to be finite. Mastouri and Zhu et al. [18] consider an infinite number of basis function in a reproducing kernel Hilbert space, and use the definitions of the inner and Kronecker products for that space. An estimate $\hat{\bm{u}}$ is obtained by solving the successive regression problems. In Stage 1, we estimate the conditional expectation $\mathbb{E}_{W|A=a,Z=z}\left[\bm{\psi}_{W}(W)\right]$ as a function of $a,z$. In Stage 2, we substitute the model (4) into the inner conditional expectation in $\mathcal{L}_{\mathrm{PV}}$,
@@
When fixed feature dictionaries are used, this two-stage regression benefits from strong theoretical guarantees [3, 18]. The use of pre-specified feature maps, limits the scope and flexibility of the method, however, especially if the treatment and proxies are high dimensional (e.g. images or text), and the causal relations are nonlinear. To overcome this, we propose to use adaptive features, expressed by neural nets, as described in the next section.
@@ 3 Deep Feature Proxy Causal Learning
In this section, we develop the DFPV algorithm, which learns adaptive features modeled by neural nets using a technique similar to Xu et al. [35]. As in Mastouri and Zhu et al. [18], we assume that we do not necessarily have access to observations from the joint distribution of $(A,Y,Z,W)$. Instead, we are given $m$ observations of $(A,Z,W)$ for Stage 1 and $n$ observations of $(A,Z,Y)$ for Stage 2. We denote the stage 1 observations by $(a_{i},z_{i},w_{i})$ and the stage 2 observations by $(\tilde{a}_{i},\tilde{z}_{i},\tilde{y}_{i})$. If observations of $(A,Y,Z,W)$ are given for both stages, we can evaluate the out-of-sample loss of Stage 1 using Stage 2 data and vice versa, and these losses can be used for hyper-parameter tuning of $\lambda_{1},\lambda_{2}$ (Appendix A). We first introduce two-stage regression with adaptive feature maps and then describe the detailed learning procedure of DFPV.
@@ 3.1 Two-stage regression with adaptive features
As in the existing work [3, 18], we learn $\mathbb{E}_{W|a,z}\left[\bm{\psi}_{\theta_{W}}(w)\right]$ in Stage 1 and $h(a,w)$ in Stage 2, but in addition to the weights $\bm{u}$ and $\bm{V}$, we also learn the parameters of the feature maps. Specifically, in Stage 1, we learn $\bm{V}$ and parameters $\theta_{A(1)},\theta_{Z}$ by minimizing the following empirical loss:
@@
Again, the loss ^$\hat{\mathcal{L}}_{2}$ is an empirical estimate of $\mathcal{L}_{2}$ in (6) with adaptive feature maps. Although the expression of ^$\hat{\mathcal{L}}_{2}$ does not explicitly contain $\theta_{W}$, it implicitly depends on $\theta_{W}$ as $(\hat{\bm{V}},\hat{\theta}_{A(1)},\hat{\theta}_{Z})$ is the solution of a minimization problem involving $\theta_{W}$. This implicit dependency makes it challenging to update $\theta_{W}$, as we cannot directly obtain its gradient. One possible solution is to use the implicit gradient method [16], but this approach might be too computationally expensive. Instead, we use the method proposed in Xu et al. [35], in which we ignore the dependency of $\theta_{W}$ on parameters $\hat{\theta}_{A(1)},\hat{\theta}_{Z}$, and compute the gradient via the closed-form solution of $\hat{V}$.
@@ 3.2 Deep Feature Proxy Variable Method
The complete procedure is presented in Algorithm 1. Note that we may use any sophisticated gradient-based learning method to optimize, such as Adam [10]. As reported in Xu et al. [35], we observe that the learning procedure is stabilized by running several gradient descent steps on the stage 1 parameters $(\theta_{A(1)},\theta_{Z})$ before updating the stage 2 features $(\theta_{A(2)},\theta_{W})$. Furthermore, we may use mini-batch updates, which sample subsets of the data at the beginning of each iteration and only use these subsamples to update the parameters.
@@ 3.3 Consistency of DFPV
In this section, we show that the DFPV method yields a consistent estimate of the bridge function. Our approach differs from the consistency result of Deaner [3] and Mastouri and Zhu et al. [18], which depend on the assumption that the true bridge function lies in a specific functional space (such as an RKHS [18]) or satisfies certain smoothness properties [3]. Since our features are adaptive, however, we instead build our result on a Rademacher complexity argument [24], which holds for a wide variety of hypothesis spaces. Furthermore, we do not assume that our hypothesis space contains the true bridge function, which makes this theory of independent interest.
@@
Assumption 3 (Completeness Assumption on Outcome-Inducing Proxy [3, 18]).
@@
We formally restate Proposition 2 in Appendix C (Theorem 3), in which the definitions of hypothesis spaces and constants can be found. From this, we can see that if we correctly identify the hypothesis space ($\kappa_{1},\kappa_{2}=0$) and have vanishing Rademacher complexities (^$\hat{\mathfrak{R}}_{S_{1}}(\mathcal{H}_{1}),\hat{\mathfrak{R}}_{S_{2}}(\mathcal{H}_{2})\to 0$ in probability), then the estimated bridge function converges in probability to the true one. Note that the Rademacher complexity of certain neural networks is known; e.g., ReLU two-layer networks are known to have the complexity of order $O(\sqrt{1/n})$ [26]. However, it might be not straightforward to derive the Rademacher complexities of the specific classes of neural networks we use, because we are employing the outer product of two neural networks in (5) and (6), whose Rademacher complexity remains a challenging open problem. The analysis in Proposition 2 is similar to Xu et al. [35], with two important distinctions: first, the target $f_{\mathrm{struct}}$ is estimated using the bridge function $h$ marginalized over one of its arguments $W$, but PCL learns the bridge function itself. Second, our result is more general and includes the misspecified case, which was not treated in the earlier work.